Solve the following DE y (4) + 5y + 4y = sin x + cos2x Good Luck
Answer #1

Colition: So given that 14% + 5yll +47 sinxt Cos20 auzulazy equation is m 4+ 5m2 +4 = 0 (m2+) cm2+4) = 0 m= 11 +21 Cof= (a coso P-I = sinx. + COS2X 1 (4+4) (02+1) (-4+1) (D249) P.I = I. Isina -I I COS27C D2+1 3 D²+4 3 we know that I simax = -X cos@26